OpenStudy (anonymous):

Solve the equations x+y=3 y=2x-15 A.(5,-5) B.(-10,-5) C.(1,2) D.(6,-3)

OpenStudy (hoblos):

x+y=3 y=2x-15 replace this equation in the first one x+2x-15=3 3x=18 x=6 y=2x-15 = 12-15=-3 (6,-3)